Circular No. 7857

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2000 CF_{105}

K. Noll and D. Stephens, Space Telescope Science Institute; W. Grundy, J. Spencer, R. Millis, and M. Buie, Lowell Observatory; D. Cruikshank, Ames Research Center, NASA; S. Tegler, Northern Arizona University; and W. Romanishin, University of Oklahoma, report the identification of a binary companion to the transneptunian object 2000 CF_{105} (MPEC 2000-F02, MPO 25770) on two exposures each with the F555W and F814W filters taken on Jan. 12.050–12.076 UT with the Hubble Space Telescope (+ WFPC2). The companion is cleanly separated in all four images at a distance of 0′′.78 ± 0′′.03 at p.a. 107^{o} ± 2^{o}; then at Δ = 41.32 AU, the two components were separated by at least 23 000 km. The spacecraft was tracking 2000 CF_{105} during the observations. Stars and galaxies moved by 1′′.1 between the first and last exposure of the sequence. The magnitudes of the primary were V = 24.25 ± 0.11 and I = 22.99 ± 0.08; those of the secondary were V = 25.12 ± 0.24 and I = 23.55 ± 0.13.

SUPERNOVA 2001ke IN ANONYMOUS GALAXY K. Z. Stanek, Center for Astrophysics (CfA); P. M. Garnavich and

S. T. Holland, University of Notre Dame; and S. Jha, R. P. Kirshner, and D. Bersier, CfA, report that R-band images of GRB 011121 (cf. GCN 1147) were obtained with the Walter Baade 6.5-m Magellan telescope on 2001 Dec. 4.3 (12.5 days after the γ-ray burst). Using point-spread-function fitting, a magnitude of R = 23.0 ± 0.1 is derived (two magnitudes brighter than expected from the power-law decline observed in the first three days after the burst; e.g., GCN 1154). Recently released Hubble Space Telescope (HST) images taken 13, 23, and 28 days after the burst (GCN 1273, 1274, 1276) confirm that the Magellan magnitude is not contaminated by host light and indicate a slow decline consistent with a supernova’s reaching peak brightness < 2 rest-frame weeks after the GRB. HST images yield a position α = 11^{h}34^{m}29^{s}.64, δ = −76^{o}01′41′′.51 (equinox J2000; uncertainty ± 0′′.2), which is 0′′.87 west and 0′′.11 north of the host-galaxy’s center. For a redshift of z = 0.36 (GCN 1152), and correcting for a large Galactic dust extinction of E(B − V) = 0.5, an absolute magnitude of M_{v} = −19.3 ± 0.2 is found for Dec. 4. This is consistent with the peak luminosity of a bright supernova, such as SN 1998bw. The reddening-corrected spectral-energy distribution from the HST photometry peaks at a rest wavelength of ≈ 500 nm (indicating a thermal source) and is inconsistent with a typical GRB power-law distribution. It is concluded that late-time observations were dominated by a supernova that was likely the origin of GRB 011121.

2002 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green